# Towards Explainable Spoofed Speech Attribution and Detection: a Probabilistic Approach for Characterizing Speech Synthesizer Components

### v2

@@ 7 Conclusion
We proposed a novel framework for probabilistic characterization of spoofed speech, aimed at enhancing explainability and interpretability in spoofing detection and attack attribution (source tracing) tasks. Our initial findings are promising. In particular, they indicate that it is feasible to extract relatively low-dimensional probabilistic embeddings without compromising predictive performance relative to the use of original raw CM embeddings. For both tasks, the reported Shapley values revealed that duration and conversion attributes play an important in spoofing detection, whereas waveform generation (vocoder) and speaker modeling are important in attribution. These encouraging results warrant further investigations both into the attribute extractors as well as the back-end classifiers.
Our findings highlight the effectiveness of the CM embedding extractor and MLPs as a foundation for generating probabilistic attribute embeddings. While this serves as a strong starting point, we see opportunities for further improvements. Moving forward, we plan to refine the framework by optimizing the embeddings directly for classifying spoofing attacks and concerned attributes, to enhance their overall effectiveness. To ensure broader applicability across diverse scenarios, we aim to evaluate the framework on the recently available MLAAD [69] and ASVspoof5 corpora [94], fostering domain and language generalization. In the present study, focused on conceptual contribution, our intentional focus has been on the ASVspoof 2019 corpus which consists of high-quality audio materials. Generalization to other types of domains and attacks will be addressed in future work. Another important direction is to evaluate the robustness of the proposed approach under real-world conditions containing nuisance variations. These include, but are not limited to, speech codec or audio compression effects and additive noise. Future efforts will focus on systematically assessing and enhancing the framework’s resilience against these factors. Overall, our work represents a modest step toward developing explainable and interpretable frameworks for robust spoofing detection and spoofing attack attribution, with promising avenues for further enhancements.
Appendix: ML estimator for the probabilistic embedding features

### v3

@@ 7 Conclusion
We proposed a novel framework for probabilistic characterization of spoofed speech, aimed at enhancing explainability and interpretability in spoofing detection and attack attribution (source tracing) tasks. Our initial findings are promising. In particular, they indicate that it is feasible to extract relatively low-dimensional probabilistic embeddings without compromising predictive performance relative to the use of original raw CM embeddings. For both tasks, the reported Shapley values revealed that waveform generator, conversion model outputs, and inputs attributes play an important in spoofing detection, whereas inputs, speaker modeling, and duration are important in attribution. These encouraging results warrant further investigations both into the attribute extractors as well as the back-end classifiers. Additionally, we observed that SSL-AASIST embeddings perform better for spoofing detection, while AASIST embeddings are more effective for spoofing attack attribution, outperforming Raw-boost variant in their respective tasks, highlighting the complementary roles of self-supervised learning and data augmentation.
Our findings highlight the effectiveness of the CM embedding extractor and MLPs as a foundation for generating probabilistic attribute embeddings. While this serves as a strong starting point, we see opportunities for further improvements. Moving forward, we plan to refine the framework by optimizing the embeddings directly for classifying spoofing attacks and concerned attributes, to enhance their overall effectiveness. To ensure broader applicability across diverse scenarios, we aim to evaluate the framework on the recently available MLAAD [69] and ASVspoof5 corpora [94], fostering domain and language generalization. In the present study, focused on conceptual contribution, our intentional focus has been on the ASVspoof 2019 corpus which consists of high-quality audio materials. Generalization to other types of domains and attacks will be addressed in future work. Another important direction is to evaluate the robustness of the proposed approach under real-world conditions containing nuisance variations. These include, but are not limited to, speech codec or audio compression effects and additive noise. Future efforts will focus on systematically assessing and enhancing the framework’s resilience against these factors. Furthermore, a promising direction is the integration of spoofing detection and spoofing attack attribution within a unified framework. This would require careful design of evaluation protocols and metrics, and will be explored in future studies. Overall, our work represents a modest step toward developing explainable and interpretable frameworks for robust spoofing detection and spoofing attack attribution, with promising avenues for further enhancements.
Appendix-1: ML estimator for the probabilistic embedding features
@@
Appendix-2: Probabilistic attributes with bonafide utterances
Recall that all our probabilistic attribute detectors have been trained using deepfake speech only. Given that they are nonetheless used as features in a task (deepfake detection) that involves bonafide samples, one may ask what is the expected behavior of these features in response to bonafide speech?
Figure 7: Distribution of probabilistic attribute values for the bonafide speech utterances.
To explore this, we used the probabilistic attribute extractors trained with the ASVspoof2019-det protocol. Figure 7 presents the output distributions from each attribute detector when applied to bonafide samples from the evaluation set. Since the attribute–value pairs are inspired by characteristics of spoofed speech generation, our hypothesis is that bonafide utterances might be assigned to the attribute values that are closest to those seen in spoofed speech. In these cases, the prediction distributions are highly polarized, with probabilities concentrated near 00 or $1$. For example, in the case of the input attribute, bonafide speech is predominantly assigned a value of $1$ for Speech and 00 for Text. Similarly, for other attributes, bonafide speech is consistently assigned to specific values: NLP for processor, None for duration and speaker, CART for conversion, LPC for output, and SpectFiltOLA for waveform, with all other values receiving near-zero probabilities.
Interestingly, the input attribute shows that bonafide speech is more closely aligned with voice conversion (VC) attacks (where input is speech) than with text-to-speech (TTS) attacks (where input is text). Moreover, whenever None is a possible attribute value, bonafide speech tends to be assigned to it. These observations suggest that the attribute–value predictions for bonafide samples tend to cluster around specific values, potentially indicating a form of implicit alignment or resemblance between certain spoofed configurations and genuine speech.
In the future, we plan to expand the set of attributes to better explain bonafide speech by incorporating characteristics related to the speech production mechanism, and integrate these with the existing attributes to enhance interpretability of the decision-making process.
Appendix-3: Impact of Unknown Attacks on the Proposed Framework
Figure 8: (a) Hamming distance and (b) confusion matrix. A16 and A19 are treated as known attacks, as their generation configurations are identical to A04 and A06, respectively. The color schemes for the Hamming distance and confusion matrix are intentionally inverted to enhance visualization, given the inverse relationship between the two. In the confusion matrix, A01–A06 correspond to spoofing attacks from the development set, while A07–A19 represent attacks from the evaluation set.
The attribution experiments in this study have been focused on the closed-set set assumption, i.e. that all test time attacks are known at the training time. It is nonetheless relevant to also analyze the response of attribution methods to unseen attacks. To this end, we used the full evaluation set (excluding bonafide samples) from the original ASVspoof2019 protocol.
Even if the unseen evaluation set attacks do not have a corresponding label in the training set, it is possible to quantify closeness (hence, confusability) of any pair of attacks through the distance of their ground-truth attribute vectors. For this purpose, we use Hamming distance of the one-hot encoded attribute vectors of known attacks (from the training and development sets) and those of unknown attacks (from the evaluation set). Hamming distance of 0 means that the pair of attacks are identical in terms of their ground-truth attributes. Using probabilistic attribute extractors trained on the training set of ASVspoof2019-attr-2, we classified evaluation set utterances—comprising unknown attacks—into the known attack classes. The resulting confusion matrix was analyzed to examine the relationship between unknown-to-known attack assignments and their corresponding Hamming distances. Our hypothesis is that unknown attacks are more likely to be assigned to known classes with smaller Hamming distances. The computed Hamming distances and the confusion matrix are presented in Figure 8.
It is evident that known attacks with a Hamming distance of zero exhibit clear correspondence, while the likelihood of correct attack assignment decreases as the Hamming distance increases. Extending this observation to unknown-to-known attack assignments, the trend generally holds, although not in all cases. This behavior suggests that improving the generalizability of both the embedding extractor and the probabilistic attribute extractor could enhance performance. In real-world scenarios—where novel spoofing attacks are continually emerging—this approach offers a principled way to assign new attacks to the most similar known classes, with the added benefit of explainability.
Acknowledgment
The work has been partially supported by the Academy of Finland (Decision No. 349605, project ”SPEECHFAKES”). The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.